OpenStudy (anonymous):

a superhero asked his nemesis how many cats she has. she answered with a riddle "seven-eighths of my cats plus eight" how many cats does the nemesis have?

OpenStudy (anonymous):

\[ c = \frac{7}{8}c+8 \]

OpenStudy (ryan123345):

Hey there! I see that you are new to openstudy. This is a great website on getting help with questions of any subject. It is a fun way of learning! I wish you good luck on openstudy and dont forget to read the http://openstudy.com/code-of-conduct so you know the rules and regulations :) Hope you have fun!

OpenStudy (anonymous):

If x is the number of cats then: x = 7/8x + 8 x - 7/8x = 8 1/8x = 8 x = 8 / (1 / 8) x = 8 * 8 x = 64 So he has 64 cats.

OpenStudy (anonymous):

how did u get 1/8?

OpenStudy (shubhamsrg):

anyways, what does the question has anything to do with "superhero" ? :/